how would you find the intercepts for the equations X+Y=56 and y/x=3/1 and where would they intersect if graphing? I am not sure I understand how to do this. Please help.
Answer by palanisamy(496)

The equation of the first line is x+y = 56 ...(1)
To find the y - intercept put x = 0
(1) => 0+y = 56
y=56
So the y intercept is ( 0, 56)
Next,
To find the x - intercept put y = 0
(1) => x+0 = 56
x=56
So the x intercept is ( 56 , 0)
The equation of the first line is y/x = 3/1 ...(2)
y=3x
This line passes through origin.
So its x intercept is (0,0)
and x intercept is (0,0)
The poit of intersection is obtained by solving (1) and (2)
I am unable to give you the graph here
For more details and clear your doubts pl contact me
kpalanisamy369@gmail.com
Put y=3x in (1)
We get x+3x=56
4x=56
x=56/4
x=14
y=3*14
y=42
So the point of intersection is ( 14,42)